I need help with this problem: It is from Week by Week Mathematics Essentials
Grade 8 Week 12
Sean paddled upstream 3 hours to camp. The next day he makes the return trip downstream in 1 hour. Sean can paddled 2 mph in still water. How fast is the current?
Thank you.

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

(2-C)3=(2+C)1
6-3C=2+C
-3C-C=2-6
-4C=-4
C=-4/-4
C=1 MPH IS THE CURRENT.
PROOF
(2-1)3=(2+1)1
1*3=3*1
3=3

Answer by stanbon(48558)

Sean paddled upstream 3 hours to camp. The next day he makes the return trip downstream in 1 hour. Sean can paddled 2 mph in still water. How fast is the current?
-----------------
Let the current speed be "c".
-------------------
Upstream DATA:
Time = 3 hrs ; Rate = 2-c mph ; Distance = rt = 3(2-c) miles
-----------------
Downstream DATA:
Time = 1 hr ; Rate = (2+c) mph ; Distance = rt = 2+c miles
-----------------
EQUATION:
distance up = distance down
6-3c = 2+c
4c = 4
c = 1 mph (speed of the current)
==============
Cheers,
Stan H.